OpenStudy (anonymous):

How would I solve this? How do I do it with Diameter instead of radius? Or can I use the diameter in place of the radius?

OpenStudy (anonymous):

I DO NOT WANT THE ANSWER! A cone has a base with a diameter of 60 cm. The height of the cone is 24 cm. What is the approximate volume of the cone? Use 3.14 for π. A. 1507.2 cubic centimeters B. 22,608 cubic centimeters C. 67,824 cubic centimeters D. 90,432 cubic centimeters

OpenStudy (anonymous):

@ganeshie8 @Hero @undeadknight26 @HelpBlahBlahBlah @campbell_st

hero (hero):

Use the volume of cone formula \[V_{\text{cone}} = \frac{\pi r^2 h}{3}\]

OpenStudy (anonymous):

I already know the formula that's not what my question was....

hero (hero):

\[r = \frac{d}{2}\]

OpenStudy (anonymous):

OK, Can you shoe me the formula now using Diameter? :)

hero (hero):

First find the radius using \(r = \dfrac{d}{2}\). Then use the Volume of Cone Formula to find the Volume of the Cone.

OpenStudy (anonymous):

So... it would be 1/3 x 3.14 x 60/2 x 24?

hero (hero):

Actually the radius is squared

OpenStudy (anonymous):

OK... 1/3 x 3.14 x 60/2^2 x 24?

hero (hero):

\((1/3)(3.14)(60/2)^2(24)\)

OpenStudy (anonymous):

Can I turn the fractions into decimals so I can put them into my calculator? How would I put that into my calculator? o_o

hero (hero):

What kind of calculator do you have that doesn't accept fractions?

hero (hero):

Use parentheses as I have demonstrated above.

hero (hero):

I typed it exactly as you would enter it in to a calculator

OpenStudy (anonymous):

an online calculator. :) just on from my computer. hold on let me see if i can put it in...

OpenStudy (anonymous):

Now I'm using Google's calculator. Can I change the fractions to decimals? It won't let me put in fractions. What would 1/3 and 60/2 be as decimals? :)

OpenStudy (anonymous):

@Hero @ganeshie8

hero (hero):

http://bit.ly/1hI9qwq